# AdvReal: Adversarial Patch Generation Framework with Application to Adversarial Safety Evaluation of Object Detection Systems

### v1

@@ Abstract
Autonomous vehicles are typical complex intelligent systems with artificial intelligence at their core. However, perception methods based on deep learning are extremely vulnerable to adversarial samples, resulting in safety accidents. How to generate effective adversarial examples in the physical world and evaluate object detection systems is a huge challenge. In this study, we propose a unified joint adversarial training framework for both 2D and 3D samples to address the challenges of intra-class diversity and environmental variations in real-world scenarios. Building upon this framework, we introduce an adversarial sample reality enhancement approach that incorporates non-rigid surface modeling and a realistic 3D matching mechanism. We compare with 5 advanced adversarial patches and evaluate their attack performance on 8 object detecotrs, including single-stage, two-stage, and transformer-based models. Extensive experiment results in digital and physical environments demonstrate that the adversarial textures generated by our method can effectively mislead the target detection model. Moreover, proposed method demonstrates excellent robustness and transferability under multi-angle attacks, varying lighting conditions, and different distance in the physical world. The demo video and code can be obtained at  https://github.com/Huangyh98/AdvReal.git.
@@
Adversarial patch , Object detection system , Reality enhanced , AI safety , Autonomous driving
@@ 1 Introduction
With the rapid advancement of artificial intelligence (AI), autonomous vehicles (AVs) have become widely adopted worldwide. Drivable area detection and object detection are the primary actors in any autonomous driving(AD) technology Zhang et al. (2022); Giri et al. (2025). Although the algorithms have achieved significant breakthroughs in performance, the frequent traffic accidents involving autonomous vehicles have raised public concerns about the safety and robustness of their perception systems. Consequently, the systematic testing and evaluation of object detection algorithms have become a prominent research focus.
@@
Currently, AD object detection systems face two primary risks: natural scenario risks and adversarial scenario risks. Natural scenario risks stem from variations in the real environment, such as changes in lighting, weather, and the behavior of traffic participants Swerdlow et al. (2024). The performance of perceptive system are typically simulated and assessed through scenario generation techniques. In contrast, adversarial scenario risks involve malicious attackers employing adversarial attacks to disrupt the normal operation of perception systems in both digital and physical domains Huang et al. (2024); Ran et al. (2025); Hu et al. (2023). Notably, physical attacks directly target objects in the real world, effectively bridging the digital and physical realms to produce a more potent effect. Such attacks can cause the autonomous driving system to fail in accurately detecting and localizing objection, which in turn may lead to hazardous driving decisions Fang et al. (2024); Bai et al. (2024).
@@
The physical adversarial attack patch can be used to effectively evaluate the security of the target detection system Zhang et al. (2021). The adversarial patch for pedestrian detection was first introduced by AdvPatch Brown et al. (2017). By optimizing and generating adversarial patches, this approach effectively evades target detectors in the physical world, although it is limited to static objects. Subsequently, researchers have proposed a variety of patch-based physical adversarial attack methods, such as AdvTshirt Xu et al. (2020), AdvTexture Hu et al. (2022), and T-SEA Huang et al. (2023). In addition, techniques like NatPatch Hu et al. (2021) and AdvCaT Hu et al. (2023) have been developed for natural-looking and camouflaged attacks. Previous studies demonstrated strong adversarial attack performance in small-scale digital tests, yet they did not account for the multifaceted effects present in real-world environments.
@@
Researchers found out that patch attacks’ threaten may pose a smaller threat than previously believed, and that the success rate observed in simple digital simulations does not necessarily translate to real-world effectiveness Hingun et al. (2023). In pedestrian detection tasks, factors such as intra-class diversity, ambient lighting, and clothing wrinkles can all diminish the performance of adversarial patches Wang et al. (2024). For instance, applying patches to three-dimensional objects may lead to geometric distortions due to perspective projection effects Mahima et al. (2024). The discrepancy between simplified adversarial training in digital environments and the complex, dynamic nature of the physical world is the primary challenge in successfully applying adversarial patches Cui et al. (2024).
@@
In this study, we proposed a reality-enhanced adversarial training framework called AdvReal. Through realistic enhancement and joint optimization of samples in both two-dimensioal (2D) and three-dimensinal (3D) spaces, more effective adversarial patch are generated for the physical world. The framework comprises three modules. The patch adversarial module refines patches in 2D and synthesizes samples that capture intra-class diversity. The realistic enhancement module processes 3D models to produce lifelike adversarial samples. And the joint optimization module concurrently refines both 2D and 3D adversarial samples to yield robust patches for complex physical world, shown in Fig. 1. The main contributions of this work are as follows:
@@
    Designed a general joint adversarial training framework for both 2D and 3D samples to overcome the challenges posed by intra-class diversity and variations in real-world environments during patch training.
@@
    Proposed an adversarial sample reality enhancement method based on non-rigid surface modeling and a realistic matching mechanism in 3D space. By simulating realistic factors, including cloth wrinkles and lighting conditions in the physical world, adversarial training and object detection evaluation can be effectively achieved using realistic 3D targets.
@@
    Numerous sets of experiments were conducted in both digital and physical environments, and we analyzed the underlying attack mechanism of the patch using visualization methods. Our quantitative and qualitative results show that the generated adversarial patch achieves a higher attack success rate, improved robustness, and superior transferability against pedestrian detectors compared to existing methods.
@@ 2.1 Adversarial Patch
Compared to digital perturbations added directly to images Xiao et al. (2025), adversarial patches present a unique capability to attack detectors in the physical world. Since these patches are installed before the camera imaging process, they pose a greater threat to autonomous driving perception system and are of greater practical significance.
As a typical sparse perturbation, the adversarial patch efficiently creates patches that can universally degrade the accuracy of the classification Williams and Li (2024). In the context of traffic sign detection, Zhou et al. investigated how adversarial patches can manipulate AVs by making traffic signs unrecognizable or misclassified, showcasing the real-world implications of such attacks Zhou et al. (2024).Adversarial patches are also widely used in face recognition, where wearable masks Liu et al. (2024) or even tiny stickers Wei et al. (2022) can be generated to mislead deep neural networks. Therefore, adversarial patches have been widely used to evaluate the security and reliability of object detection system.
@@ 2.2 Evade Person Detectors with Adversarial Patch
Humans are among the most critical traffic participants. Using adversarial patches to attack pedestrian detection system provides an effective means to evaluate their susceptibility to adversarial risks. Unlike objects such as traffic signs, which exhibit minimal intra-class variability, humans present significant intra-class diversity, making the detection task more challenging Abed et al. (2024).
A common approach involves optimizing adversarial patches by placing them on the body and minimizing the maximum bounding-box probability associated with human detection. AdvPatch was the first adversarial patch attack method to deceive pedestrian detectors by optimizing detection loss Thys et al. (2019). Similarly, TOG employs attack-specific gradients based on adversarial targets to cause detectors to either miss or misclassify pedestrians Chow et al. (2020). However, TOG does not account for applications in the physical world. Furthermore, a GAN-based generative strategy restricts the appearance of generated patches to ensure that they look natural and inconspicuous Guesmi et al. (2024). NatPatch achieves this by sampling optimal images from GANs Hu et al. (2021), resulting in adversarial patches that are both visually natural and highly effective in attacks. All of these methods rely on real images of diverse individuals, overlaying adversarial patches within bounding boxes in a digital environment in 2D picture.
@@ 2.3.1 High level of intra-class variety
Pedestrians exhibit a high degree of intra-class variability, characterized by significant differences in appearance, posture, and dynamic motion. To address these variations, the SOTA pedestrian detection models leverage extensive training data sets and advanced feature extraction techniques to improve the robustness of the algorithms Ren et al. (2016a); Redmon (2018); Carion et al. (2020). Consequently, devising effective and generalized adversarial attack strategies requires exceptional adaptability and robustness. Simen Thys et al. were among the first to explore such attacks on targets with high intra-class variability, utilizing real images of diverse individuals Thys et al. (2019). Building on this, Zhou et al. proposed the PosePatch framework, a patch adaptation network for adversarial patch synthesis guided by perspective transformations and estimated human poses Zhou et al. (2025). These advancements underscore the importance of leveraging abundant intra-class training samples to significantly enhance the robustness and effectiveness of adversarial patches.
@@ 2.3.2 Multi-angle attacks
Traditional superposition training ensures that pedestrian detectors perform effectively when facing the patch but lacks robustness to varied angles or occlusions. To address the challenges posed by multi-angle perception attacks, AdvTexture introduces a scalable adversarial texture generation method based on ring cropping Hu et al. (2022). This approach can be applied to clothing of arbitrary shapes, allowing individuals wearing these garments to evade detection by human detectors from various angles. Furthermore, Arman Maesumi et al. proposed a framework for adversarial training using 3D modeling, which optimizes 3D adversarial cloaks through variations in poses, spatial positions, and camera angles Maesumi et al. (2021). By conducting multi-angle adversarial training, the effectiveness of patches in multi-detection angle scenarios can be significantly improved Li et al. (2025); Hu et al. (2023).
@@ 2.3.3 Realistic factors
Accurately capturing all real-world variations through digital images of simulated patches for adversarial training and evaluation remains challenging Hingun et al. (2023). Modeling non-rigid object surfaces, such as clothing, during training and utilizing them for adversarial optimization poses a considerable challenge. For instance, Adv-Tshirt employs video frames of a person wearing a T-shirt with a checkerboard pattern as training data, mapping the patch onto a distorted checkerboard for training Xu et al. (2020). AdvCaT proposes a non-rigid deformation method based on TopoProj to simulate wrinkles on human clothing Hu et al. (2023). The inclusion of adversarial training methods that account for non-rigid wrinkles enhances the robustness of patches in real-world physical environments. In addition, Traditional patch superposition methods assume that patches are square and axis aligned and can be placed anywhere within the image. However, random rotation, cropping, and noise do not fully simulate real-world variations in lighting, noise, and posture. REAP proposes an improved method that incorporates geometric and lighting transformations to enable large-scale performance evaluation of patches, particularly traffic signal signs Hingun et al. (2023). In summary, the reality augmentation method can effectively improve the robustness of the adversarial patch for realistic challenges.
@@ 3.1 Problem Definition
In this study, we conduct a joint attack on 2D high-level intra-class varity data and 3D augmented reality data. The ultimate goal is to effectively reduce the accuracy of adversarial patches successfully recognized by human detectors in the real world. Specifically, let $P(X,Y)$ represent the joint distribution of pedestrian data, where $X\subseteq\mathbb{R}^{d}$ is the input image space and $Y\subseteq\{1,\ldots,C\}$ is the label space. A pretrained pedestrian detector $f_{h}:[-1,1]^{d}\to[0,1]^{C}$ maps the input $x\in X$ to a probability distribution over $C$ classes, with the predicted label $\hat{y}=\arg\max_{c}f_{c}(x)$. Given samples $\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}\sim P(X,Y)$, the goal of adversarial attacks is to construct a perturbed input $x^{\text{adv}}=x+\delta$, where $\|\delta\|_{p}\leq\epsilon$, such that the detector $f(x^{\text{adv}})$ misclassifies $x^{\text{adv}}$ into a label $\hat{y}\neq y$.
The optimization problem is then formulated as Eq. 1:
[…]
where $\mathcal{L}(f(x_{i}+\delta),y_{i})$ measures the discrepancy between the detector’s prediction on the perturbed input $x_{i}+\delta$ and the true label $y_{i}$. The regularization term $\mathcal{R}(\delta)$ ensuring the smoothness of the adversarial patch, thereby reducing visual artifacts, and improving the generalization and transferability of the patch by preventing over-fitting. The coefficient $\lambda$ balances the adversarial loss and the regularization term, ensuring that the adversarial patch is both effective in disrupting the detector’s performance and robust enough to generalize across diverse real-world scenarios and detection architectures.

### v2

# AdvReal: Physical Adversarial Patch Generation Framework for Security Evaluation of Object Detection Systems

@@ Abstract
Autonomous vehicles are typical complex intelligent systems with artificial intelligence at their core. However, perception methods based on deep learning are extremely vulnerable to adversarial samples, resulting in security accidents. How to generate effective adversarial examples in the physical world and evaluate object detection systems is a huge challenge. In this study, we propose a unified joint adversarial training framework for both 2D and 3D domains, which simultaneously optimizes texture maps in 2D image and 3D mesh spaces to better address intra-class diversity and real-world environmental variations. The framework includes a novel realistic enhanced adversarial module, with time-space and relighting mapping pipeline that adjusts illumination consistency between adversarial patches and target garments under varied viewpoints. Building upon this, we develop a realism enhancement mechanism that incorporates non-rigid deformation modeling and texture remapping to ensure alignment with the human body’s non-rigid surfaces in 3D scenes. Extensive experiment results in digital and physical environments demonstrate that the adversarial textures generated by our method can effectively mislead the target detection model. Specifically, our method achieves an average attack success rate (ASR) of 70.13% on YOLOv12 in physical scenarios, significantly outperforming existing methods such as T-SEA (21.65%) and AdvTexture (19.70%). Moreover, the proposed method maintains stable ASR across multiple viewpoints and distances, with an average attack success rate exceeding 90% under both frontal and oblique views at a distance of 4 meters. This confirms the method’s strong robustness and transferability under multi-angle attacks, varying lighting conditions, and real-world distances. The demo video and code can be obtained at  https://github.com/Huangyh98/AdvReal.git.
@@
AI Security , Object detection system , Adversarial attack , Augmented reality , Autonomous driving perception
@@ 1 Introduction
With the rapid advancement of artificial intelligence (AI), autonomous vehicles (AVs) have become widely adopted worldwide. Drivable area detection and object detection are the primary components in any autonomous driving (AD) technology [30, 49]. Although these algorithms have achieved significant breakthroughs in performance, the frequent traffic accidents involving autonomous vehicles have raised public concerns about the security and robustness of their perception systems. Consequently, the systematic testing and evaluation of object detection algorithms have become a prominent research focus.
@@
Currently, AD object detection systems face two primary risks: natural scenario risks and adversarial scenario risks. Natural scenario risks stem from variations in the real environment, such as changes in lighting, weather, and the behavior of traffic participants [32]. The performance of perception systems is typically simulated and assessed through scenario generation techniques. In contrast, adversarial scenario risks involve malicious attackers employing adversarial attacks to disrupt the normal operation of perception systems in both digital and physical domains [33, 48, 19]. Notably, physical attacks directly target objects in the real world, effectively bridging the digital and physical realms and resulting in more impactful consequences. Such attacks can cause the autonomous driving system to fail in accurately detecting and localizing objects, which in turn may lead to hazardous driving decisions [50, 31].
@@
Physical adversarial patches can be used to effectively evaluate the security of object detection systems [51]. The adversarial patch for pedestrian detection was first introduced by AdvPatch [1]. By optimizing and generating adversarial patches, this approach effectively evades target detectors in the physical world, although it is limited to static objects. Subsequently, researchers have proposed a variety of patch-based physical adversarial attack methods, such as AdvTshirt [14], AdvTexture [12], and T-SEA [27]. In addition, techniques like NatPatch [16] and AdvCaT [19] have been developed for natural-looking and camouflaged attacks. While previous studies demonstrated strong attack performance in constrained digital settings, they often neglected the complex environmental factors encountered in real-world applications.
@@
Researchers have found that the threat posed by patch-based attacks may be smaller than previously believed, and that the success rate observed in simple digital simulations does not necessarily translate to real-world effectiveness [20]. In pedestrian detection tasks, factors such as intra-class diversity, ambient lighting, and clothing wrinkles can all diminish the performance of adversarial patches [34]. For instance, applying patches to three-dimensional objects may lead to geometric distortions due to perspective projection effects [36]. The discrepancy between simplified adversarial training in digital environments and the complex, dynamic nature of the physical world is one of the primary challenges hindering the real-world deployment of adversarial patches [35].
@@
In this study, we proposed a reality-enhanced adversarial training framework called AdvReal. Through realistic enhancement and joint optimization of samples in both two-dimensional (2D) and three-dimensional (3D) spaces, more effective adversarial patches are generated for the physical world. The framework comprises three modules. The patch adversarial module refines patches in 2D and synthesizes samples that capture intra-class diversity. The realistic enhancement module processes 3D models to produce lifelike adversarial samples. And the joint optimization module concurrently refines both 2D and 3D adversarial samples to yield robust patches for complex physical world scenarios, shown in Fig. 1. Our goal is to generate adversarial patches to evaluate and expose vulnerabilities in perception systems deployed in AVs, especially pedestrian detection modules, under real-world environmental conditions. The main contributions of this work are as follows:
@@
    Designed a general joint adversarial training framework for both 2D and 3D samples to overcome the challenges posed by intra-class diversity and variations in real-world environments during patch training. The framework supports the simulation of realistic adversarial threats against AV perception models by capturing differences in viewpoint, distance, and body appearance common in urban driving scenarios.
@@
    Proposed an adversarial sample reality enhancement method based on non-rigid surfaces modeling and a realistic matching mechanism in 3D space. By incorporating realistic elements such as cloth deformation and varying illumination, the method better reflects how adversarial patches would appear on pedestrians in actual AV perception settings.
@@
    Numerous sets of experiments were conducted in both digital and physical environments, and we analyzed the underlying attack mechanism of the patch using visualization methods. Our quantitative and qualitative results show that the generated adversarial patch achieves a higher attack success rate, improved robustness, and superior transferability against pedestrian detectors compared to existing methods. The results offer actionable insights for improving detector architectures in AV systems and contribute to the broader goal of AI security evaluation in autonomous driving.
@@ 2.1 Adversarial Patch
Understanding the behavior of adversarial attackers has become a growing focus in AI security research. Recent works have proposed systematic taxonomies to describe attacker characteristics, including access level (e.g., glass-box, gray-box, closed-box), behavior characteristics (e.g., digital and physical attack), attack intention (e.g., evasion, impersonation, sabotage, et al.), and attack capability (e.g., attack success rate, transferability, robustness, et al.) [54]. In adversarial attack tasks, a glass-box (or white-box) setting assumes the attacker has full access to the victim model’s parameters and gradients, whereas a closed-box (or black-box) setting assumes no internal knowledge of the model is available, and the attacker can only interact with the detector via input-output queries. The frameworks help researchers to analyze attack feasibility, threat severity, and defense applicability across different deployment scenarios. In safety-critical systems such as autonomous driving, distinguishing attacker intent and capability is especially important, as it directly impacts how risks should be assessed and mitigated [60]. The background provides a foundation for studying concrete physical-world attacks such as adversarial patches.
Compared to digital perturbations added directly to images [52], adversarial patches present a unique capability to attack detectors in the physical world. Since these patches are installed before the camera imaging process, they pose a greater threat to autonomous driving perception systems and are of greater practical significance.
As a typical sparse perturbation, an adversarial patch can effectively degrade classification performance across various inputs [2]. In the context of traffic sign detection, Zhou et al. investigated how adversarial patches can manipulate AVs by making traffic signs unrecognizable or misclassified, showcasing the real-world implications of such attacks [3]. Adversarial patches are also widely used in face recognition, where wearable masks [5] or even tiny stickers [4] can be generated to mislead deep neural networks. Therefore, adversarial patches have been widely used to evaluate the security and reliability of object detection systems.
@@ 2.2 Evade Person Detectors with Adversarial Patch
Humans are among the most critical traffic participants in autonomous driving environments. Using adversarial patches to attack pedestrian perception modules in AV systems provides an effective means to evaluate their susceptibility to adversarial risks. Unlike objects such as traffic signs, which exhibit minimal intra-class variability, humans present significant intra-class diversity, making the detection task more challenging [7] and increasing the difficulty of ensuring robust and reliable performance under adversarial conditions.
A common approach involves optimizing adversarial patches by placing them on the human body and minimizing the maximum bounding-box probability associated with pedestrian detection. AdvPatch was the first adversarial patch method designed to fool pedestrian detectors by directly optimizing detection loss [8]. Similarly, TOG employs attack-specific gradients derived from adversarial targets to cause detectors to either miss or misclassify pedestrians [15]. However, TOG does not account for deployment in physical-world settings. Furthermore, a GAN-based generative strategy has been introduced to constrain the appearance of generated patches, ensuring they remain natural and inconspicuous [17]. NatPatch follows a similar principle by sampling optimal images from GANs [16], resulting in patches that are both visually realistic and adversarially effective. While the above methods offer promising results in static image-based evaluations, they fall short of addressing the full complexity of AV perception systems, which must operate under dynamic viewpoints, environmental variations, and real-time physical constraints. As such, their ability to evaluate adversarial robustness in practical autonomous driving scenarios remains limited.
@@ 2.3.1 High level of intra-class variety
Pedestrians exhibit a high degree of intra-class variability, characterized by significant differences in appearance, posture, and dynamic motion. In autonomous driving perception systems, the variability poses major challenges for consistent and reliable pedestrian detection under real-world conditions. To address these variations, the SOTA pedestrian detection models leverage extensive training datasets and advanced feature extraction techniques to improve algorithmic robustness [9, 10, 11]. Consequently, devising effective and generalized adversarial attack strategies requires exceptional adaptability and resilience. Simen Thys et al. were among the first to explore such attacks on targets with high intra-class variability, utilizing real images of diverse individuals [8]. Building on this, Zhou et al. proposed the PosePatch framework, a patch adaptation network for adversarial patch synthesis guided by perspective transformations and estimated human poses [6]. The efforts reveal the necessity of accounting for real-world appearance diversity when evaluating the AI security of pedestrian detection modules in AVs.
@@ 2.3.2 Multi-angle attacks
Traditional superposition training ensures that pedestrian detectors perform effectively when facing the patch but lacks robustness to varied angles or occlusions. In the context of AVs, where onboard cameras observe pedestrians from dynamic and shifting perspectives, adversarial patches must be effective across a wide range of viewing angles to pose a realistic threat. To address these challenges, AdvTexture introduces a scalable adversarial texture generation method based on ring cropping [12]. This approach can be applied to clothing of arbitrary shapes, allowing individuals wearing these garments to evade detection by human detectors from various angles. Furthermore, Arman Maesumi et al. proposed a framework for adversarial training using 3D modeling, which optimizes 3D adversarial cloaks through variations in poses, spatial positions, and camera angles [13]. By conducting multi-angle adversarial training, the effectiveness of patches in multi-detection angle scenarios can be significantly improved [38, 19]. Such techniques are critical for testing the vulnerability of real-world pedestrian detection systems deployed in autonomous vehicles.
@@ 2.3.3 Realistic factors
Accurately capturing all real-world variations through digital images of simulated patches for adversarial training and evaluation remains challenging [20]. Modeling non-rigid object surfaces, such as clothing, during training and utilizing them for adversarial optimization poses a considerable challenge. For instance, Adv-Tshirt employs video frames of a person wearing a T-shirt with a checkerboard pattern as training data, mapping the patch onto a distorted checkerboard for training [14]. AdvCaT proposes a non-rigid deformation method based on TopoProj to simulate wrinkles on human clothing [19]. Integrating such non-rigid modeling into the patch optimization process significantly improves robustness under real-world physical constraints. Traditional patch-based methods often assume that patches are square, axis-aligned, and freely placed within the image. However, random rotation, cropping, and additive noise do not fully simulate complex environmental variations such as illumination shifts, pose changes, and background clutter. REAP proposes an improved method that incorporates geometric and lighting transformations to enable large-scale performance evaluation of patches, particularly traffic signal signs [20]. To ensure safety and reliability in autonomous driving applications, it is essential to include physically realistic transformations in patch training pipelines to evaluate adversarial robustness under dynamic AV scenarios.
@@ 3.1 Problem Definition
In this study, we perform a joint adversarial attack on both 2D high-level intra-class variability data and 3D augmented reality data. The ultimate objective is to effectively reduce the detection accuracy of adversarial patches recognized by human detectors in real-world settings. Specifically, let $P(X,Y)$ denote the joint distribution of pedestrian data, where $X\subseteq\mathbb{R}^{d}$ is the input image space and $Y\subseteq\{1,\ldots,C\}$ is the label space. A pretrained pedestrian detector $f_{h}:[-1,1]^{d}\to[0,1]^{C}$ maps an input $x\in X$ to a probability distribution over $C$ classes, with the predicted label $\hat{y}=\arg\max_{c}f_{c}(x)$. Given a sample set $\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}\sim P(X,Y)$, the goal of the adversarial attack is to construct a perturbed input $x^{\text{adv}}=x+\delta$, where $\|\delta\|_{p}\leq\epsilon$, such that the detector $f(x^{\text{adv}})$ misclassifies $x^{\text{adv}}$ into a label $\hat{y}\neq y$.
The corresponding optimization problem is formulated in Eq. 1:
$\delta^{*}=\underset{\delta}{\arg\min}\;\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(f_{h}(x_{i}+\delta),y_{i})+\lambda\mathcal{R}(\delta),$ (1)
where $\mathcal{L}(f(x_{i}+\delta),y_{i})$ measures the discrepancy between the detector’s prediction on the perturbed input $x_{i}+\delta$ and the true label $y_{i}$. The regularization term $\mathcal{R}(\delta)$ enforces the smoothness of the adversarial patch, thereby reducing visual artifacts and improving both its generalization and transferability by preventing overfitting. The coefficient $\lambda$ balances the adversarial loss and the regularization term, ensuring that the adversarial patch is not only effective in misleading the detector but also robust across diverse real-world scenarios and detection architectures.